OpenStudy (anonymous):

-6x+9=-21 plz can some give me this answer I been working on this all DAY!!!!!!!

OpenStudy (phi):

did you (as a first step ) add -9 to both sides ? what did you get ?

OpenStudy (anonymous):

I DOT KNOW WAT THAT MEANS

OpenStudy (anonymous):

DONT

OpenStudy (anonymous):

srry caps

OpenStudy (phi):

it means you write -9 on both sides of the = sign, like this -6x+9 -9 =-21 -9 now simplify

OpenStudy (anonymous):

im guessing

OpenStudy (phi):

are you saying you can't simplify 9 - 9 on the left side ? or you can't do -21 - 9 on the right side?

OpenStudy (anonymous):

-21-9

OpenStudy (anonymous):

srry I have been working on math all day and I have about 5 more hours of it

OpenStudy (anonymous):

i just want to get done

OpenStudy (phi):

what is 9 - 9 ?

OpenStudy (anonymous):

0

OpenStudy (phi):

what is -21 - 9 ?

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

-30

OpenStudy (phi):

yes. When you have time, you should practice doing it in your head. you started with -6x + 9 = -21 you add -9 to both sides -6x+9 -9 =-21 -9 you then got -6x + 0 = -30 on the left side -6x + 0 is just -6x so you have -6x = -30 divide both sides by -6 can you do that ?

OpenStudy (anonymous):

how do you divide -6x by -6

OpenStudy (anonymous):

plz can u plz give me this answer cuz i got many more questions

OpenStudy (phi):

-6 divided by -6 is 1 \[ \frac{-6 x }{-6}= 1x= x \]

OpenStudy (phi):

now do -30/-6 on the right side

OpenStudy (anonymous):

5

OpenStudy (phi):

so -6x = -30 divide both sides by -6, and you get x = 5 that is the answer

OpenStudy (anonymous):

thx so so so so so so so so so so so MUCH!!!!!!!!!!!!!!!!!